Twelve goats were set loose in an island and their population growth can be approximated by the function p(t)=[65(t-1)/t+7] How many goats will there by 6 years
1
Expert's answer
2021-10-12T08:04:38-0400

in the initial year,

"p(1)=\dfrac{65(1-1)}{1+7}= 0"


in the second year,

"p(2)=\dfrac{65(2-1)}{2+7}= \dfrac{65}{9} = 7.22"


in the third year,

"p(3)=\dfrac{65(3-1)}{3+7}= \dfrac{130}{9}= 14.44"


in the fourth year,

"p(4)=\dfrac{65(4-1)}{4+7}=\dfrac{195}{11} = 17.73"


in the fifth year,

"p(5)=\dfrac{65(5-1)}{5+7}=\dfrac{260}{12}= 21.67"


in the sixth year,

"p(6)=\dfrac{65(6-1)}{6+7}=\dfrac{325}{13}=25"


The total number of goats after 6 years will be "0+7.22+14.44+17.73+21.67+25= 86.06\approx 86\textsf{ goats}"